TVM

The Maybe Pay Life Insurance Co. is trying to sell you an investment policy that will pay you and your heirs $25,000 per year forever. If the required return on this investment is 7.2 percent, how much will you pay for the policy?
Answer #1

The formula for the present value of a perpetuity payment is as follows

PV = C/r

C: payment received

r: requried rate of returns

 

In this case:

The payment received= $ 25,000

Required rate of return= 7.2%

PV= 25000/7.2%= $347,222

=> I need to pay $347,222 for the policy


answered by: trang trần